# Investigation of magnetic excitations and charge order in a van der Waals ferromagnet Fe5GeTe2

### v2

@@ 2.3 Absence of charge order in bulk Fe5GeTe2
Overall, our REXS and XRD studies report the presence of a commensurate (1/3 1/3 ) Bragg peak with a 3+1.5 out-of-plane modulation, no temperature dependence down to 7 K, and no resonance effect. Based on these observations, we propose these peaks to be of structural origin rather than electronic [23, 24, 25, 34]. This conclusion is further corroborated by the comparable peak intensity with respect to main structural Bragg peaks. Furthermore, we underline that the observed 3+1.5 out-of-plane periodicity is not compatible with the space group proposed for Fe5GeTe2, instead, it infers doubling of structural unit cell along the -axis.
@@
To reconcile the discrepancy between our observations and the ones reported in the literature [33, 17], we should acknowledge intrinsic differences connected with the techniques involved. STM and ARPES mostly probe the surface, which in certain materials can significantly differ from the bulk [35, 36, 37, 38]. Electron diffraction instead probes the bulk, but the required sample preparation methods - such as focused ion beam and grinding - might introduce permanent micro-structural changes affecting the material response [39, 40]. Finally, we cannot exclude a sample to sample variation in determining the ultimate material properties. Fe5GeTe2 single crystals are known to have stacking faults and different samples (especially grown by different groups) might even have slightly different Fe concentration.
3 Conclusion and outlook
@@
Our work sheds light on the magnetic excitations and electronic order in the room temperature vdW ferromagnet Fe5GeTe2. Using high-resolution RIXS at the Fe edge, we reveal that Fe5GeTe2 hosts dual character of magnetic excitations composed of a magnon and a continuum, similarly to the more extensively characterized Fe3GeTe2, and prove that the dual character of magnetic excitation driven by orbital-selective Mott transition is common to ternary Fe-Ge-Te vdW materials. The out-of-plane modulation of the continuum intensity reveals the presence of a strong out-of-plane inter-slab exchange path, where the interacting magnetic units are the neighboring Fe-Ge-Te slabs. We conclude that Fe5GeTe2 behaves like a 3D magnet, and we discuss this finding in connection to its , the highest among Fe-Ge-Te vdW materials. Additionally, employing XRD and REXS, we explored the bulk character, the electronic origin and the propagation vector of the previously reported charge order with (1/3 1/3) in-plane components. The absence of temperature and resonance effects, and the commensurability with the crystal structure lead us to conclude a structural origin for the observed Bragg peaks. The identified (1/3 1/3 1.5) propagation vector requires doubling of the crystallographic unit cell.
Acknowledgments
This work was supported as part of Programmable Quantum Materials, an Energy Frontier Research Center funded by the US Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0019443. Development of Fe5GeTe2 crystals (AFM) was supported by the U. S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. This research used beamlines 2-ID and 4-ID of NSLS-II, a US DOE Office of Science User Facility operated for the DOE Office of Science by Brookhaven National Laboratory under contract no. DE-SC0012704. This research used resources of the Advanced Photon Source, a U.S. Department of Energy (DOE) Office of Science user facility operated for the DOE Office of Science by Argonne National Laboratory under Contract No. DE-AC02-06CH11357.
Author Contributions
VBi, VBh, CM, AM, DB, and XX conceived the research plan. AFM grew and characterized the bulk single crystalline Fe5GeTe2. VBi, VBh, TK, JL, YG, SF, JP carried out the RIXS measurements. VBh, TK, CM, CN, JF performed the XRD and REXS experiments. TD and DJR realized the Raman measurements. VBh and VBi analyzed and interpreted the data, with contributions from all coauthors. VBh and VBi wrote the manuscript. All authors commented on the text.
@@ Methods
The Raman measurements were taken using a LabRAM HR Evolution Raman microscope from Horiba Scientific using 100 0.9NA Olympus air objective. Spectra were recorded using an 1800 gr/mm holographic grating and Synapse II front-illuminated EMCCD.

### v3

@@ 2.2 Origin of the magnetic continuum
Overall, our REXS and XRD studies report the presence of a commensurate (1/3 1/3 ) Bragg peak with a 3+1.5 out-of-plane modulation, no temperature dependence down to 7 K, and no resonance effect. Based on these observations, we propose these peaks originate from the crystal structure rather than charge order [25, 26, 27, 46]. This conclusion is further corroborated by the comparable peak intensity with respect to main structural Bragg peaks. Furthermore, we underline that the observed 3+1.5 out-of-plane periodicity is not compatible with the space group proposed for Fe4.75GeTe2, instead, it infers doubling of structural unit cell along the -axis and reduced crystal symmetry. One direct consequence of reduced lattice symmetry would be a more complex network of magnetic interactions, which might be crucial to interpret the various kinks in the magnetization versus temperature diagram [4].
@@
To reconcile the discrepancy between our observations and the ones reported in the literature [45, 17], we should acknowledge intrinsic differences connected with the techniques involved. STM and ARPES mostly probe the surface, which in certain materials can significantly differ from the bulk [47, 48, 49, 50]. Electron diffraction instead probes the bulk, but the required sample preparation methods - such as focused ion beam and grinding - might introduce permanent micro-structural changes affecting the material response [51, 52]. Finally, we cannot exclude a sample to sample variation in determining the ultimate material properties. Fe4.75GeTe2 single crystals are known to have stacking faults and different samples (especially grown by different groups) might even have slightly different Fe concentration. To advance the understanding of this complex system, it would be helpful to perform on the same sample both surface sensitive (STM, ARPES) and bulk sensitive (XRD, REXS) investigations.
@@
In summary, our work sheds light on the magnetic excitations and electronic order in the room temperature vdW ferromagnet Fe4.75GeTe2. Using high-resolution RIXS at the Fe edge, we reveal that Fe4.75GeTe2 hosts dual character of magnetic excitations composed of a magnon and a continuum, similarly to the more extensively characterized Fe2.72GeTe2, and prove that the dual character of magnetic excitation, known to be driven by orbital selective Mott transition, is common to ternary Fe-Ge-Te vdW materials. The out-of-plane modulation of the continuum intensity reveals the presence of a strong out-of-plane inter-slab exchange path, where the interacting magnetic units are the neighboring Fe-Ge-Te slabs. We conclude that Fe4.75GeTe2 behaves like a 3D magnet, and we discuss this finding in connection to its , the highest among Fe-Ge-Te vdW materials. Additionally, employing XRD and REXS, we explored the bulk character, the electronic origin and the propagation vector of the previously reported charge order with (1/3 1/3) in-plane components. The absence of temperature and resonance effects, and the commensurability with the crystal structure led us to conclude that, the observed Bragg peaks do not originate from charge order but rather crystal structure with reduced symmetry. The presence of the (1/3 1/3 3+1.5) peaks suggests a doubling of the unit cell along the c-axis, indicating a subtle change in the stacking sequence or atomic positions that is not captured by the space group.
@@ Methods
We performed room-temperature Raman scattering ( = 6328 Å) on a Fe4.75GeTe2 flake exfoliated on SiO2/Si and oriented as -axis normal using a LabRAM HR Evolution Raman microscope from Horiba Scientific using 100 0.9NA Olympus air objective. Spectra were recorded using an 1800 gr/mm holographic grating and Synapse II front-illuminated EMCCD (Supplementary Figure 1).
Data availability
Data supporting the findings of this study are available from the corresponding author on reasonable request.
Acknowledgments
This work was supported as part of Programmable Quantum Materials, an Energy Frontier Research Center funded by the US Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0019443. Development of Fe4.75GeTe2 crystals (AFM) was supported by the U. S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. This research used beamlines 2-ID and 4-ID of NSLS-II, a US DOE Office of Science User Facility operated for the DOE Office of Science by Brookhaven National Laboratory under contract no. DE-SC0012704. This research used resources of the Advanced Photon Source, a U.S. Department of Energy (DOE) Office of Science user facility operated for the DOE Office of Science by Argonne National Laboratory under Contract No. DE-AC02-06CH11357.
Author Contributions
VBi, VBh, CM, AM, DB, and XX conceived the research plan. AFM grew and characterized the bulk single crystalline Fe4.75GeTe2. VBi, VBh, TK, JL, YG, SF, JP carried out the RIXS measurements. VBh, TK, CM, CN, JF performed the XRD and REXS experiments. TD and DJR realized the Raman measurements. VBh and VBi analyzed and interpreted the data, with contributions from all coauthors. VBh and VBi wrote the manuscript. All authors commented on the text.
3 Competing Interests
The authors declare no competing interests.